OpenStudy (anonymous):

What is a counterexample of the following conjecture? Conjecture: Any number that is divisible by 3 is also divisible by 6. A.27 B.36. c.23. d.18

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Any number that is divisible by both 2 and 3 are divisible by 6. 3^2 = 9 9/3 = 3 9/6 = 1.5

OpenStudy (anonymous):

Could you explain to me what is a counterexample? Are they looking for a number that can or can't?

OpenStudy (anonymous):

Counter example is an example that proves it false

OpenStudy (anonymous):

Oh, just saw you have options

OpenStudy (anonymous):

Divide each by 3 and 6, and see which one isn't a whole number